OpenStudy (anonymous):

the manager of a coffee shop has one type of coffee that sells for $7 per pound and another type that sells for $11 per pound. The manager wishes to mix 60 pounds of the $11 coffee to get a mixture that will sell for $8 per pound. How many pounds of the $7 coffee should be used?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

help

OpenStudy (anonymous):

let he mixes x pound of $7 coffe then by the problem we have 11*60 +7x =8*(x+60) or 660+7x=8x+480 or x=180